OpenStudy (faiqraees):

An electric fish is capable of generating a voltage of 220V. This drives a pulse of 15A lasting for 2ms. The fish produces 200 pulses in one second. What is the average power output?

OpenStudy (faiqraees):

At first glance I thought we just have to multiply 220 with 15 using the equation P=VI. But that doesn't give the answer. So at another glance I was convinced that 200 is also supposed to be multiplied because in one second 200 such "power" is produced. But that still doesn't gives the answer. Can anyone explain what am I missing?

OpenStudy (irishboy123):

Energy of 1 pulse = \(VI \Delta t = 220 \times 15 \times 2/1000\) Energy in 1 sec = \(200 \times 220 \times 15 \times 2/1000\) which is also the power as its \(P = \dfrac{\Delta E}{\Delta t} = \dfrac{E}{1}\) does that give you the answer you need?

OpenStudy (faiqraees):

Why energy in 1 sec = 2ms * Energy of 1 pulse?

OpenStudy (irishboy123):

you get 200 pulses in 1 sec, right?

OpenStudy (irishboy123):

so you get the energy of 200 pulses in 1 second. and \(P = \dfrac{E}{t}\), and here t = 1

OpenStudy (faiqraees):

and why is 2ms multiplied here

OpenStudy (irishboy123):

to get the Energy of a single pulse. E = P * t

OpenStudy (irishboy123):

P = IV so E = IV * t

OpenStudy (faiqraees):

Oh got it thanks